An urn contains four red marbles and five blue marbles. What is the probability of selecting at random, without replacement, two blue marbles?
Answer by longjonsilver(2297)

i have a lesson on probability on this site. Go and read it.

Anyway, to answer you, without much explanation:

P(BB) = P(first B) AND P(second B)
P(BB) = 5/9 * 4/8

Once we select the first blue and do not replace it, we are left with the urn holding 4 red and 4 blue... hence the second fraction in my calculation.

P(BB) = 5/9 * 1/2
P(BB) = 5/18

jon.